# Highly edge-connected subgraphs

### v1

@@ Abstract
    There exists a multi-graph with minimum degree 2k12k-1 without a $(k+1)$-edge-connected subgraph: Consider the graph $G$ on vertex set $\{1,\dots,2(k+1)\}$ with k1k-1 parallel edges between each two consecutive vertices $i$ and $i+1$, one additional edge between $1$ and each of the $k$ vertices $\{2,\dots,k+1\}$, and between $2(k+1)$ and each of the $k$ vertices $\{k+2,\dots,2k+1\}$. Clearly, all vertices have degree 2k12k-1. Between the two vertex sets $\{1,\dots,k+1\}$ and $\{k+2,\dots,2k+2\}$ there are only k1k-1 edges. Hence, if $G^{\prime}$ is a $(k+1)$-connected subgraph of $G$, then either $V(G^{\prime})\subseteq\{1,\dots,k+1\}$ or $V(G^{\prime})\subseteq\{k+2,\dots,2k+2\}$. In both cases, there is a vertex of degree at most $k$, $\max V(G^{\prime})$ or $\min V(G^{\prime})$, respectively. This is a contradiction.
@@
    A similar construction can be used for simple graphs. Together with ideas from Theorem On highly edge-connected subgraphs, one can show that for $k\leq 5$, a minimum degree 2k12k-1 is still best possible: One can build graphs of minimum degree 2k12k-1 without a $(k+1)$-connected subgraph. Because the construction allows no generalization, we won’t show it here. Instead, we now consider large $k$.
We now prove that, for simple graphs and for $k\to\infty$, already a minimum degree $\big{(}1+o(1)\big{)}k$ is sufficient for a $(k+1)$-edge-connected subgraph.
Let $\alpha\in\big{(}0,\frac{1}{2}\big{)}$ be fixed. All of the following definitions will depend on $k$, which will not be noticeable in the notations. Let $m=m(k)\in[0,k]$ not necessarily an integer, which we will choose later.
Let $G$ be a graph and $v\in V(G)$ of degree $d_{G}(v)$. We define a penalty
$m_{G}(v)\leavevmode\nobreak\ :=\leavevmode\nobreak\ 0\lor\big{(}k+m-d_{G}(v)\big{)}\land m\leavevmode\nobreak\ \in[0,m]\,.$
@@
$\sum_{j=1}^{l}\alpha j^{\alpha-1}\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{j=1}^{l}\int_{j-1}^{j}\alpha j^{\alpha-1}dx\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{j=1}^{l}\int_{j-1}^{j}\alpha x^{\alpha-1}dx\leavevmode\nobreak\ =\leavevmode\nobreak\ \int_{0}^{l}\alpha x^{\alpha-1}dx\leavevmode\nobreak\ =\leavevmode\nobreak\ l^{\alpha}\,.$
@@
$\sum_{j=1}^{l}\alpha j^{\alpha-1}\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{j=1}^{l}\int_{j}^{j+1}\alpha j^{\alpha-1}dx\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \sum_{j=1}^{l}\int_{j}^{j+1}\alpha x^{\alpha-1}dx\leavevmode\nobreak\ =\leavevmode\nobreak\ \int_{1}^{l+1}\alpha x^{\alpha-1}dx\leavevmode\nobreak\ \geq\leavevmode\nobreak\ l^{\alpha}-1\,.$
@@
$M(G)\leavevmode\nobreak\ \geq\leavevmode\nobreak\ M(G-w)-k^{\alpha}\,.$
@@
$M(G-w)\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{j=1}^{v(G)-1}\alpha j^{\alpha-1}m_{G-w}(v_{j})\,.$
For every $j\in\{1,\dots,v(G)-1\}$, we have $d_{G}(v_{j})-d_{G-w}(v_{j})\in\{0,1\}$ as $G$ is simple. Since the function $x\mapsto 0\lor\big{(}(k+m)-x\big{)}\land m$ is monotone decreasing and non-expanding, this yields $\Delta_{j}:=m_{G-w}(v_{j})-m_{G}(v_{j})\in[0,1]$ and
$\sum_{j=1}^{v(G)-1}\Delta_{j}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{j=1}^{v(G)-1}d_{G}(v_{j})-d_{G-w}(v_{j})\leavevmode\nobreak\ =\leavevmode\nobreak\ d_{G}(w)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ k\,.$
@@
$M(G)\leavevmode\nobreak$ $\geq\leavevmode\nobreak\ \sum_{j=1}^{v(G)-1}\alpha j^{\alpha-1}m_{G}(v_{v(G)-1})+\alpha v(G)^{\alpha-1}m_{G}(w)$
$>\leavevmode\nobreak\ \sum_{j=1}^{v(G)-1}\alpha j^{\alpha-1}\big{(}m_{G-w}(v_{j})-\Delta_{j}\big{)}\leavevmode\nobreak\ =\leavevmode\nobreak\ M(G-w)-\sum_{j=1}^{v(G)-1}\alpha j^{\alpha-1}\Delta_{j}\,.$
Since $\sum_{j=1}^{v(G)-1}\Delta_{j}\leq k$ and the weights $\alpha j^{\alpha-1}$ are decreasing, the weighted sum over the $\Delta_{j}$ is maximized if $\Delta_{j}=1$ for $j\leq k$ and $\Delta_{j}=0$ otherwise. That is,
$M(G)\leavevmode\nobreak\ \geq\leavevmode\nobreak\ M(G-w)-\sum_{j=1}^{k}\alpha j^{\alpha-1}\leavevmode\nobreak\ \geq\leavevmode\nobreak\ M(G-w)-k^{\alpha}\,$,
using the upper bound of the first $k$ weights. $\Box$
@@
$M(G)\leavevmode\nobreak\ \geq\leavevmode\nobreak\ 2^{\alpha-1}\big{(}M(A)+M(B)\big{)}-m\Big{(}\frac{4k}{m}\Big{)}^{\alpha}\,.$
@@
$M(A)\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{j=1}^{v(A)}\alpha j^{\alpha-1}m_{A}(a_{j})\leavevmode\nobreak\ \leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ \leavevmode\nobreak\ M(B)\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{j=1}^{v(B)}\alpha j^{\alpha-1}m_{B}(b_{j})\,.$
Clearly $d_{A}(a_{j})=d_{G-S}(a_{j})$ and hence $m_{A}(a_{j})=m_{G-S}(a_{j})$. Similarly $m_{B}(b_{j})=m_{G-S}(b_{j})$. We may assume by symmetry that $v(A)\leq v(B)$. Consider the following list of V(GS)=V(G)V(G-S)=V(G):
$a_{1},b_{1},a_{2},b_{2},\dots,a_{v(A)},b_{v(A)},\dots,b_{v(B)}\,.$
We obtain
$M(G-S)\leavevmode\nobreak$ $\geq\leavevmode\nobreak\ \sum_{j=1}^{v(A)}\alpha(2j-1)^{\alpha-1}m_{A}(a_{j})+\sum_{j=1}^{v(A)}\alpha(2j)^{\alpha-1}m_{B}(b_{j})+\hskip-10.00002pt\sum_{j=v(A)+1}^{v(B)}\hskip-10.00002pt\alpha(v(A)+j)^{\alpha-1}m_{B}(b_{j})$
$\geq\leavevmode\nobreak\ \sum_{j=1}^{v(A)}\alpha(2j)^{\alpha-1}m_{A}(a_{j})+\sum_{j=1}^{v(B)}\alpha(2j)^{\alpha-1}m_{B}(b_{j})\leavevmode\nobreak\ =\leavevmode\nobreak 2^{\alpha-1}M(A)+2^{\alpha-1}M(B)\,.$
@@
$M(G-S)\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{j=1}^{v(G)}\alpha j^{\alpha-1}m_{G}(v_{j})\,.$
For every $j\in\{1,\dots,v(G)\}$, we have $d_{G}(v_{j})\geq d_{G-S}(v_{j})$, and hence $\Delta_{j}:=m_{G-S}(v_{j})-m_{G}(v_{j})\in[0,m]$. It is
$\sum_{j=1}^{v(G)}\Delta_{j}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{j=1}^{v(G)}d_{G}(v_{j})-d_{G-S}(v_{j})\leavevmode\nobreak\ =\leavevmode\nobreak\ 2|S|\leavevmode\nobreak\ \leq\leavevmode\nobreak\ 2k\,.$
@@
$M(G)\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \sum_{j=1}^{v(G)}\alpha j^{\alpha-1}\big{(}m_{G-S}(v_{j})-\Delta_{j}\big{)}\leavevmode\nobreak\ =\leavevmode\nobreak\ M(G-S)-\sum_{j=1}^{v(G)}\alpha j^{\alpha-1}\Delta_{j}\,.$
Since $\sum_{j=1}^{v(G)}\Delta_{j}\leq 2k$ and the weights $\alpha j^{\alpha-1}$ are decreasing, we obtain an upper bound of the weighted sum over the $\Delta_{j}$ with $\Delta_{j}=m$ for $j\leq\big{\lceil}\frac{2k}{m}\big{\rceil}$ and $\Delta_{j}=0$ otherwise. That is,
$M(G)\leavevmode\nobreak\ \geq\leavevmode\nobreak\ M(G-S)-\sum_{j=1}^{\big{\lceil}\frac{2k}{m}\big{\rceil}}\alpha j^{\alpha-1}m\leavevmode\nobreak\ \geq\leavevmode\nobreak\ 2^{\alpha-1}\big{(}M(A)+M(B)\big{)}-m\Big{(}\frac{4k}{m}\Big{)}^{\alpha}\,$,
using the upper bound of the first $\big{\lceil}\frac{2k}{m}\big{\rceil}\leq\frac{4k}{m}$ weights. $\Box$
@@
Proof.  Let $\alpha\in\big{(}0,\frac{1}{2}\big{)}$$m=k^{\frac{1}{2}+\alpha}\leq k$  and  $\mu=\frac{1}{2}m\sqrt{k}^{\alpha}$.
@@
A sequence of vertices $w_{1},\dots,w_{l}\in V(G)$ is called detachable from $G$, if for each $j\in\{1,\dots,l\}$, $w_{j}$ has degree at most $k$ in $G-w_{1}-\dots-w_{j-1}$. We set
$z(G)\leavevmode\nobreak\ :=\leavevmode\nobreak\ \max\big{\{}l\leavevmode\nobreak\ :\leavevmode\nobreak\ w_{1},\dots,w_{l}\in V(G)\text{ is detachable from }G\big{\}}\,.$
@@
Then, there is a detachable sequence $w_{1},\dots,w_{\big{\lceil}\sqrt{k}\big{\rceil}}\in V(G)$. For each $j\in\{1,\dots,l\}$$d_{G}(w_{j})\leq k+(j-1)$ and hence $m_{G}(w_{j})\geq m-(j-1)\geq m-(\big{\lceil}\sqrt{k}\big{\rceil}-1)\geq m-\sqrt{k}\geq\frac{2}{3}m$,for $k$ large enough. This implies, again for $k$ large enough,
$M(G)\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \sum_{j=1}^{\big{\lceil}\sqrt{k}\big{\rceil}}\alpha j^{\alpha-1}m_{G}(v_{j})\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \frac{2}{3}m\sum_{j=1}^{\big{\lceil}\sqrt{k}\big{\rceil}}\alpha j^{\alpha-1}\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \frac{2}{3}m(\sqrt{k}^{\alpha}-1)\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \frac{1}{2}m\sqrt{k}^{\alpha}\leavevmode\nobreak\ =\leavevmode\nobreak\ \mu\,.$
@@
Since $d_{G^{\prime}}(a)>k$ for all $a\in V(A)$, we have
$(v(A)-1)k\leavevmode\nobreak\ =\leavevmode\nobreak\ v(A)k-k\leavevmode\nobreak\ <\sum_{a\in V(A)}d_{G^{\prime}}(a)\leavevmode\nobreak\ -|S|\leavevmode\nobreak\ =\sum_{a\in V(A)}d_{A}(a)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ v(A)(v(A)-1)\,$,
@@
$M(G^{\prime})\leavevmode\nobreak\ \geq\leavevmode\nobreak\ 2^{\alpha-1}\big{(}M(A)+M(B)\big{)}-m\Big{(}\frac{4k}{m}\Big{)}^{\alpha}\leavevmode\nobreak\ =\leavevmode\nobreak\ 2^{\alpha}\mu-m\,o(\sqrt{k})^{\alpha}\leavevmode\nobreak\ =\leavevmode\nobreak\ 2^{\alpha}\mu-o(\mu)\,.$
Finally, applying Lemma On highly edge-connected subgraphsa on each of the vertices $w_{1},\dots,w_{z(G)}$ yields
$M(G)\leavevmode\nobreak\ \geq\leavevmode\nobreak\ M(G-w_{1}-\dots-w_{z(G)})-z(G)k^{\alpha}\leavevmode\nobreak\ \geq\leavevmode\nobreak\ M(G^{\prime})-k^{\frac{1}{2}}k^{\alpha}\leavevmode\nobreak\ =\leavevmode\nobreak\ 2^{\alpha}\mu-o(\mu)\,.$
Hence, for $k$ large enough, $M(G)\geq\mu$.
@@
For every constant $C$, there exists a $k\in\mathbb{N}$ and a graph with minimum degree at least $k+C\sqrt{k}$ without a $(k+1)$-edge-connected subgraph.
@@
    $|A_{j}|\leavevmode\nobreak\ =\leavevmode\nobreak\ 2^{j}(C\sqrt{k}+1)$.
@@
Define $G_{0}$ as a graph on $k+C\sqrt{k}+1$ vertices that contains an anticlique $A_{0}$ of $C\sqrt{k}+1$ vertices, a clique $V(G_{0})\setminus A_{0}$ on $k$ vertices and all possible edges between $A_{0}$ and $V(G_{0})\setminus A_{0}$. The vertices in $A_{0}$ have degree $k$ and those in $V(G_{0})\setminus A_{0}$ have degree $k+C\sqrt{k}$. A $(k+1)$-connected subgraph of $G_{0}$ cannot contain any vertex from $A_{0}$, but also $G_{0}-A_{0}$ has to few vertices to contain a $(k+1)$-connected subgraph.
@@
$|A_{j}|\leavevmode\nobreak\ =\leavevmode\nobreak\ 2^{j}(C\sqrt{k}+1)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ 2^{j+1}C\sqrt{k}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ 2^{2C^{2}}C\sqrt{k}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ k\,.$
We build a graph $H_{j}$ from $G_{j}$ by adding an anticlique $B_{j}$ on $\frac{1}{2C}\sqrt{k}\in\mathbb{N}$ vertices. From each vertex in $B_{j}$, we add one edge to every vertex in $A_{j}$ and $k-|A_{j}|$ more edges to other vertices from $G_{j}$. In $H_{j}$, the vertices in $B_{j}$ have degree $k$, the vertices in $A_{j}$ now have degree at least $k+\frac{1}{2C}(j+1)\sqrt{k}$ and all other vertices still have degree at least $k+C\sqrt{k}$. A $(k+1)$-connected subgraph of $H_{j}$ cannot contain any vertex from $B_{j}$, but also $G_{j}=H_{j}-B_{j}$ does not contain one by the assumption.
@@
The Theorems from the present paper have some applications for other issues. We want to mention decomposition problems of the following type, which were first studied by Thomassen [6]: What is the lowest number $k=k(s,t)$ such that every kk-(edge-)connected graph $G$ admits a partition of $V(G)$ into $A$ and $B$ such that $G[A]$ is ss-(edge-)connected and $G[B]$ is tt-(edge-)connected?
Corollary.   Let $s,t\geq 1$.
@@
    Let $G$ be an $(s+t+1)$-connected graph of minimum degree at least 3s+3t13s+3t-1. Then $V(G)$ has a partition into two sets $A$ of size >2s>2s and $B$ of size >2t>2t such that $G[A]$ is $(s+1)$-connected and $G[B]$ is $(t+1)$-connected.
@@
    Let $G$ be a triangle-free and $(s+t+1)$-connected graph of minimum degree at least $2s+2t$. Then $V(G)$ has a partition into two sets $A$ and $B$ such that $G[A]$ is $(s+1)$-connected and $G[B]$ is $(t+1)$-connected.
@@
    There is a function $f(r)=\big{(}1+o(1)\big{)}r$ such that the following holds: Let $G$ be an $(s+t+1)$-edge-connected graph of minimum degree at least $f(s+t)$. Then $V(G)$ has a partition into two sets $A$ and $B$ such that $G[A]$ is $(s+1)$-edge-connected and $G[B]$ is $(t+1)$-edge-connected.
Proof.  The approach is due to Thomassen [6].

### v2

@@ Abstract
    There exists a multigraph with minimum degree 2k12k-1 without a $(k+1)$-edge-connected subgraph: Consider the graph $G$ on the vertex set $\{1,\dots,2(k+1)\}$ with k1k-1 parallel edges between each two consecutive vertices $i$ and $i+1$, one additional edge between $1$ and each of the $k$ vertices $\{2,\dots,k+1\}$, and between $2(k+1)$ and each of the $k$ vertices $\{k+2,\dots,2k+1\}$. Clearly, all vertices have degree 2k12k-1. Between the two vertex sets $\{1,\dots,k+1\}$ and $\{k+2,\dots,2k+2\}$ there are only k1k-1 edges. Hence, if $G^{\prime}$ is a $(k+1)$-edge-connected subgraph of $G$, then either $V(G^{\prime})\subseteq\{1,\dots,k+1\}$ or $V(G^{\prime})\subseteq\{k+2,\dots,2k+2\}$. In both cases, there is a vertex of degree at most $k$, $\max V(G^{\prime})$ or $\min V(G^{\prime})$, respectively. This is a contradiction.
@@
    A similar construction can be used for (simple) graphs. Together with ideas from Theorem On highly edge-connected subgraphs, one can show that for $k\leq 5$, a minimum degree $2k$ is still best possible: One can build graphs of minimum degree 2k12k-1 without a $(k+1)$-edge-connected subgraph. Because the construction allows no generalization, we will not show it here. Instead, we now consider large $k$.
We now prove that, for (simple) graphs and for $k\to\infty$, already a minimum degree $\big(1+o(1)\big)k$ is sufficient for a $(k+1)$-edge-connected subgraph.
Let $\alpha\in\big(0,\frac{1}{2}\big)$ be fixed. All of the following definitions will depend on $k$, which will not be reflected in the notations. Let $m=m(k)\in[0,k]$ not necessarily an integer, which we will choose later.
Let $G$ be a graph and $v\in V(G)$ of degree $d_{G}(v)$. We define a delimited penalty
$m_{G}(v)\penalty 10000\ :=\penalty 10000\ \min\Big\{\max\big\{k+m-d_{G}(v),\,0\big\},\,m\Big\}\penalty 10000\ \in[0,m]\,.$
@@
$\sum_{j=1}^{l}\alpha j^{\alpha-1}\penalty 10000\ =\penalty 10000\ \sum_{j=1}^{l}\int_{j-1}^{j}\alpha j^{\alpha-1}dx\penalty 10000\ \leq\penalty 10000\ \sum_{j=1}^{l}\int_{j-1}^{j}\alpha x^{\alpha-1}dx\penalty 10000\ =\penalty 10000\ \int_{0}^{l}\alpha x^{\alpha-1}dx\penalty 10000\ =\penalty 10000\ l^{\alpha}\,.$
@@
$\sum_{j=1}^{l}\alpha j^{\alpha-1}\penalty 10000\ =\penalty 10000\ \sum_{j=1}^{l}\int_{j}^{j+1}\alpha j^{\alpha-1}dx\penalty 10000\ \geq\penalty 10000\ \sum_{j=1}^{l}\int_{j}^{j+1}\alpha x^{\alpha-1}dx\penalty 10000\ =\penalty 10000\ \int_{1}^{l+1}\alpha x^{\alpha-1}dx\penalty 10000\ \geq\penalty 10000\ l^{\alpha}-1\,.$
@@
$M(G)\penalty 10000\ \geq\penalty 10000\ M(G-w)-k^{\alpha}\,.$
@@
$M(G-w)\penalty 10000\ =\penalty 10000\ \sum_{j=1}^{v(G)-1}\alpha j^{\alpha-1}m_{G-w}(v_{j})\,.$
For every $j\in\{1,\dots,v(G)-1\}$, we have $d_{G}(v_{j})-d_{G-w}(v_{j})\in\{0,1\}$ as $G$ is simple. Since the function $x\mapsto\min\big\{\max\{k+m-x,\,0\},\,m\big\}$ is monotone decreasing and non-expanding, we have $m_{G-w}(v_{j})-m_{G}(v_{j})\leq d_{G}(v_{j})-d_{G-w}(v_{j})$. Hence $\Delta_{j}:=m_{G-w}(v_{j})-m_{G}(v_{j})\in[0,1]$ and
$\sum_{j=1}^{v(G)-1}\Delta_{j}\penalty 10000\ \leq\penalty 10000\ \sum_{j=1}^{v(G)-1}d_{G}(v_{j})-d_{G-w}(v_{j})\penalty 10000\ =\penalty 10000\ d_{G}(w)\penalty 10000\ \leq\penalty 10000\ k\,.$
@@
[…]
$>\penalty 10000\ \sum_{j=1}^{v(G)-1}\alpha j^{\alpha-1}\big(m_{G-w}(v_{j})-\Delta_{j}\big)\penalty 10000\ =\penalty 10000\ M(G-w)-\sum_{j=1}^{v(G)-1}\alpha j^{\alpha-1}\Delta_{j}\,.$
Since $\sum_{j=1}^{v(G)-1}\Delta_{j}\leq k$ and the weights $\alpha j^{\alpha-1}$ are decreasing, the weighted sum over the $\Delta_{j}\in[0,1]$ is maximized if $\Delta_{j}=1$ for $j\leq k$ and $\Delta_{j}=0$ otherwise. That is,
$M(G)\penalty 10000\ \geq\penalty 10000\ M(G-w)-\sum_{j=1}^{k}\alpha j^{\alpha-1}\penalty 10000\ \geq\penalty 10000\ M(G-w)-k^{\alpha}\,$,
using the upper bound on the sum of the first $k$ weights. $\Box$
@@
$M(G)\penalty 10000\ \geq\penalty 10000\ 2^{\alpha-1}\big(M(A)+M(B)\big)-m\Big(\frac{4k}{m}\Big)^{\alpha}\,.$
@@
$M(A)\penalty 10000\ =\penalty 10000\ \sum_{j=1}^{v(A)}\alpha j^{\alpha-1}m_{A}(a_{j})\penalty 10000\ \penalty 10000\ \text{ and }\penalty 10000\ \penalty 10000\ M(B)\penalty 10000\ =\penalty 10000\ \sum_{j=1}^{v(B)}\alpha j^{\alpha-1}m_{B}(b_{j})\,.$
Clearly $d_{A}(a_{j})=d_{G-S}(a_{j})$ and hence $m_{A}(a_{j})=m_{G-S}(a_{j})$. Similarly $m_{B}(b_{j})=m_{G-S}(b_{j})$. We may assume by symmetry that $v(A)\leq v(B)$. By considering the list
$a_{1},b_{1},a_{2},b_{2},\dots,a_{v(A)},b_{v(A)},b_{v(A)+1},\dots,b_{v(B)}$
of all vertices of V(GS)=V(G)V(G-S)=V(G), we obtain
[…]
@@
$M(G-S)\penalty 10000\ =\penalty 10000\ \sum_{j=1}^{v(G)}\alpha j^{\alpha-1}m_{G-S}(v_{j})\,.$
We have $d_{G}(v_{j})\geq d_{G-S}(v_{j})$ and hence $\Delta_{j}:=m_{G-S}(v_{j})-m_{G}(v_{j})\in[0,m]$ for every $j\in\{1,\dots,v(G)\}$. It is
$\sum_{j=1}^{v(G)}\Delta_{j}\penalty 10000\ \leq\penalty 10000\ \sum_{j=1}^{v(G)}d_{G}(v_{j})-d_{G-S}(v_{j})\penalty 10000\ =\penalty 10000\ 2|S|\penalty 10000\ \leq\penalty 10000\ 2k\,.$
@@
$M(G)\penalty 10000\ \geq\penalty 10000\ \sum_{j=1}^{v(G)}\alpha j^{\alpha-1}m_{G}(v_{j})\penalty 10000\ =\penalty 10000\ \sum_{j=1}^{v(G)}\alpha j^{\alpha-1}\big(m_{G-S}(v_{j})-\Delta_{j}\big)\penalty 10000\ =\penalty 10000\ M(G-S)-\sum_{j=1}^{v(G)}\alpha j^{\alpha-1}\Delta_{j}\,.$
Since $\sum_{j=1}^{v(G)}\Delta_{j}\leq 2k$ and the weights $\alpha j^{\alpha-1}$ are decreasing, we obtain an upper bound of the weighted sum over the $\Delta_{j}\in[0,m]$ by setting $\Delta_{j}=m$ for $j\leq\big\lceil\frac{2k}{m}\big\rceil$ and $\Delta_{j}=0$ otherwise. That is,
$M(G)\penalty 10000\ \geq\penalty 10000\ M(G-S)-\sum_{j=1}^{\big\lceil\frac{2k}{m}\big\rceil}\alpha j^{\alpha-1}m\penalty 10000\ \geq\penalty 10000\ 2^{\alpha-1}\big(M(A)+M(B)\big)-m\Big(\frac{4k}{m}\Big)^{\alpha}\,$,
using the upper bound on the sum of the first $\big\lceil\frac{2k}{m}\big\rceil\leq\frac{4k}{m}$ weights. $\Box$
@@
Proof. Let $\alpha\in\big(0,\frac{1}{2}\big)$, $m=k^{\frac{1}{2}+\alpha}\leq k$ and $\mu=\frac{1}{2}m\sqrt{k}^{\alpha}$.
@@
A sequence of pairwise distinct vertices $w_{1},\dots,w_{l}\in V(G)$ is called detachable from $G$, if for each $j\in\{1,\dots,l\}$, $w_{j}$ has degree at most $k$ in $G-w_{1}-\dots-w_{j-1}$. We set
$z(G)\penalty 10000\ :=\penalty 10000\ \max\big\{l\penalty 10000\ :\penalty 10000\ w_{1},\dots,w_{l}\in V(G)\text{ is detachable from }G\big\}\,.$
@@
Then, there is a detachable sequence $w_{1},\dots,w_{\big\lceil\sqrt{k}\big\rceil}\in V(G)$. For each $j\in\{1,\dots,\big\lceil\sqrt{k}\big\rceil\}$$d_{G}(w_{j})\leq k+(j-1)$ and hence $m_{G}(w_{j})\geq m-(j-1)\geq m-(\big\lceil\sqrt{k}\big\rceil-1)\geq m-\sqrt{k}\geq\frac{2}{3}m$,for $k$ large enough. Using the lower bound on the sum of the first weights, this implies, again for $k$ large enough,
$M(G)\penalty 10000\ \geq\penalty 10000\ \sum_{j=1}^{\big\lceil\sqrt{k}\big\rceil}\alpha j^{\alpha-1}m_{G}(v_{j})\penalty 10000\ \geq\penalty 10000\ \frac{2}{3}m\sum_{j=1}^{\big\lceil\sqrt{k}\big\rceil}\alpha j^{\alpha-1}\penalty 10000\ \geq\penalty 10000\ \frac{2}{3}m(\sqrt{k}^{\alpha}-1)\penalty 10000\ \geq\penalty 10000\ \frac{1}{2}m\sqrt{k}^{\alpha}\penalty 10000\ =\penalty 10000\ \mu\,.$
@@
Since $d_{G^{\prime}}(a)>k$ for all $a\in V(A)$, we have $v(A)>1$ and
$(v(A)-1)k\penalty 10000\ =\penalty 10000\ v(A)k-k\penalty 10000\ <\sum_{a\in V(A)}d_{G^{\prime}}(a)\penalty 10000\ -|S|\penalty 10000\ =\sum_{a\in V(A)}d_{A}(a)\penalty 10000\ \leq\penalty 10000\ v(A)(v(A)-1)\,$,
@@
$M(G^{\prime})\penalty 10000\ \geq\penalty 10000\ 2^{\alpha-1}\big(M(A)+M(B)\big)-m\Big(\frac{4k}{m}\Big)^{\alpha}\penalty 10000\ =\penalty 10000\ 2^{\alpha}\mu-m\Big(\frac{4k}{m}\Big)^{\alpha}\,.$
Finally we apply Lemma On highly edge-connected subgraphsa to get from $G$ to $G^{\prime}$ by removing the vertices $w_{1},\dots,w_{z(G)}$ one after another
$M(G)\penalty 10000\ \geq\penalty 10000\ M(G^{\prime})-z(G)k^{\alpha}\penalty 10000\ \geq\penalty 10000\ 2^{\alpha}\mu-m\Big(\frac{4k}{m}\Big)^{\alpha}-k^{\frac{1}{2}}k^{\alpha}\,.$
For $k\to\infty$, both $m\big(\frac{4k}{m}\big)^{\alpha}=mo(k)^{\alpha}=o(\mu)$ and $k^{\frac{1}{2}}k^{\alpha}=o(m)k^{\alpha}=o(\mu)$. Hence, for $k$ large enough, the sum of both numbers is smaller than $(2^{\alpha}-1)\mu$, that is, $M(G)\geq\mu$.
@@
For every constant $C$, there exists a $k\in\mathbb{N}$ and a (simple) graph with minimum degree at least $k+C\sqrt{k}$ without a $(k+1)$-edge-connected subgraph.
@@
    $|A_{j}|\penalty 10000\ =\penalty 10000\ 2^{j}(C\sqrt{k}+1)$.
@@
Define $G_{0}$ as a graph on $k+C\sqrt{k}+1$ vertices that contains an anticlique $A_{0}$ of $C\sqrt{k}+1$ vertices, a clique $V(G_{0})\setminus A_{0}$ on $k$ vertices and all possible edges between $A_{0}$ and $V(G_{0})\setminus A_{0}$. The vertices in $A_{0}$ have degree $k$ and those in $V(G_{0})\setminus A_{0}$ have degree $k+C\sqrt{k}$. A $(k+1)$-edge-connected subgraph of $G_{0}$ cannot contain any vertex from $A_{0}$, but also $G_{0}-A_{0}$ has too few vertices to contain a $(k+1)$-edge-connected subgraph.
@@
$|A_{j}|\penalty 10000\ =\penalty 10000\ 2^{j}(C\sqrt{k}+1)\penalty 10000\ \leq\penalty 10000\ 2^{j+1}C\sqrt{k}\penalty 10000\ \leq\penalty 10000\ 2^{2C^{2}}C\sqrt{k}\penalty 10000\ \leq\penalty 10000\ k\,.$
We build a graph $H_{j}$ from $G_{j}$ by adding an anticlique $B_{j}$ on $\frac{1}{2C}\sqrt{k}\in\mathbb{N}$ vertices. From each vertex in $B_{j}$, we add one edge to every vertex in $A_{j}$ and $k-|A_{j}|$ more edges to other vertices from $G_{j}$. In $H_{j}$, the vertices in $B_{j}$ have degree $k$, the vertices in $A_{j}$ now have degree at least $k+\frac{1}{2C}(j+1)\sqrt{k}$ and all other vertices still have degree at least $k+C\sqrt{k}$. A $(k+1)$-edge-connected subgraph of $H_{j}$ cannot contain any vertex from $B_{j}$, but also $G_{j}=H_{j}-B_{j}$ does not contain one by the assumption.
@@
The Theorems from the present paper have some applications to other issues. We want to mention decomposition problems of the following type, which were first studied by Thomassen [7]: What is the lowest number $k=k(s,t)$ such that every kk-(edge-)connected graph $G$ admits a partition of $V(G)$ into $A$ and $B$ such that $G[A]$ is ss-(edge-)connected and $G[B]$ is tt-(edge-)connected?
Corollary. Let $k_{1},k_{2}\geq 1$.
@@
    Let $G$ be a $(k_{1}+k_{2}+1)$-connected graph of minimum degree at least $3(k_{1}+k_{2})-1$. Then $V(G)$ has a partition into $C_{1}$ and $C_{2}$ such that, for $i\in\{1,2\}$, $|C_{i}|>2k_{i}$ and $G[C_{i}]$ is $(k_{i}+1)$-connected.
@@
    Let $G$ be a triangle-free $(k_{1}+k_{2}+1)$-connected graph of minimum degree at least $2(k_{1}+k_{2})$. Then $V(G)$ has a partition into $C_{1}$ and $C_{2}$ such that, for $i\in\{1,2\}$, $G[C_{i}]$ is $(k_{i}+1)$-connected (and hence $|C_{i}|\geq 2(k+1)$).
@@
    Let $G$ be a $(k_{1}+k_{2}+1)$-edge-connected graph of average degree at least $2(k_{1}+k_{2})+2$. Then $V(G)$ has a partition into $C_{1}$ and $C_{2}$ such that, for $i\in\{1,2\}$, $|C_{i}|>2k_{i}$ and $G[C_{i}]$ is $(k_{i}+1)$-edge-connected.
  4. (iv)
    There is a function $f(r)=\big(1+o(1)\big)r$ such that the following holds: Let $G$ be a $(k_{1}+k_{2}+1)$-edge-connected graph of minimum degree at least $f(k_{1}+k_{2})$. Then $V(G)$ has a partition into $C_{1}$ and $C_{2}$ such that, for $i\in\{1,2\}$, $G[C_{i}]$ is $(k_{i}+1)$-edge-connected.
Proof. The approach is due to Thomassen [7] and Stiebitz [8].